Please I need help! My question is:
With a head wind , an airplane travels 480 miles in two hours. Returning with a tail wind it takes 1 hour and 30 minutes. Find the airplane's speed and the wind speed.

Answer by richwmiller(17219)

(a+w)*1.5=480
a+w=480/1.5
a+w=320
(a-w)*2=480
a-w=240
a+w=320
add
2a=560
a=280
w=40